Help please!!
-3^(2-x) =-81

Answer by stanbon(75887)

-3^(2-x) =-81
---------
Multiply both sides by -1 to get:
3^(2-x) = 3^4
2-x = 4
x = -2
===============
Cheers,
Stan H.